### v1

# Collimation of Fast Radio Burster 20201124A; Repeaters vs. Apparent Non-Repeaters

Abstract
The recent report of a period in the active repeating Fast Radio Burster 20201124A and measurement of its spindown rate place bounds on the solid angle of its emission on the basis of energetics. The bound depends on the (unknown) efficiency of conversion of rotational energy to coherent radio emission and implies a lower bound on the Lorentz factor of the radiating charges. Bursts may be emitted along the magnetic dipole axis, in repeaters aligned with the rotational axis and the line of sight but misaligned in apparent non-repeaters. This may explain the difficulty of finding periodicity in repeaters and the low duty cycle of apparent non-repeaters.
Magnetars
1 Introduction
Fast Radio Burst (FRB) energetics remain enigmatic. The report (, 2025) of a 1.7 s period during two observations, 37 days apart, of the very active repeating FRB 20201124A implied that at least this FRB is produced by a rotating magnetic neutron star, and the similarity of the two reported periods supported their reality. The statistical significance of these detections was recently challenged (Gazith & Zackay, 2025). Zhang et al. (2025) reported a correlation between the the times of the claimed periodicities and changes in Rotation Measure, which may support the reality of the periodicities. The present paper, a preliminary version of which was published (Katz, 2025) before that challenge, discusses inferences that may be drawn if such periodicity is confirmed, either for FRB 20201124A or for some other repeating FRB.
A third epoch period determination is required to exclude (or demonstrate) that the change in period is the result of orbital motion. If the small difference (about one part in 1000) of the measured periods is attributed to spindown, the spindown age is 44 years, implying a very young object. The inferred surface magnetic field is $\sim 10^{15}$ Gauss; such objects are called magnetars. Ockham’s razor then suggests that all repeating FRB are produced by magnetars.
This Research Note assumes that the bursts of FRB 20201124A are powered by spindown rather than magnetostatic energy, that FRB have the same mechanism as radio pulsars (PSR) rather than soft gamma repeaters (SGR): FRB resemble PSR in emitting brief bursts of coherent radio radiation, sometimes with bandwidth-duration product approaching the uncertainty principle limit (Katz, 2024a), rather than the smoothly varying thermal gamma-rays of SGR.
2 Energetics
If there is no energy reservoir intermediate between rotational energy and radiated energy, an assumption justified by the $\sim 30\,\mu$s relaxation time of a neutron star magnetosphere, then comparing the spindown power inferred from the period derivative to the observed flux during bursts sets an upper limit to the solid angle into which the bursts are emitted. This implies a lower bound on the Lorentz factors of the radiating charge bunches. It also implies an upper bound on the angular divergence of the field lines on which the radiating charges move, and hence on the size of the radiating region. This last bound is unrelated to the angular width of radiation of individual charge bunches.
The spindown power of FRB 20201124A
[…]
where $I_{45}$ is the neutron star’s moment of inertia in units of $10^{45}$ g-cm2. The numerical value is obtained from the observed (, 2025) spin frequency $\omega$ and derivative $\dot{\omega}$.
FRB 20201124A is identified (, 2021) with a galaxy at $z=0.0979$, a distance of 400 Mpc. At this distance, the power radiated by a source with flux density $F_{Jy}$ in Janskys into a solid angle $\Omega$ over the 500 MHz observing bandwidth is
[…]
Defining $\epsilon$ as the efficiency of converting spindown power to radiation in the observing bandwidth, Eqs. 1 and 2 yield
$\Omega\approx 4\pi\left({I_{45}\over F_{Jy}}\right)\left({\epsilon\over 0.01}\right)\times 2\times 10^{-7}\ \text{sterad}.$ (3)
Known radio pulsars have $10^{-6}\lesssim\epsilon\lesssim 10^{-2}$. The brightest bursts of FRB 20201124A have $F_{Jy}$ up to 10 (, 2025). Taking $I_{45}=1$ and $F_{Jy}=10$,
$\Omega\approx 2.5\times 10^{-7}\left({\epsilon\over 0.01}\right)\ \text{sterad}.$ (4)
For a conical beam of half-angle $\theta$
$\theta\approx\sqrt{\Omega\over\pi}\approx 3\times 10^{-4}\sqrt{\epsilon\over 0.01}\ \text{rad}.$ (5)
This sets a lower bound on the Lorentz factor of the radiating charges
$\Gamma\gtrsim\theta^{-1}\sim 3\times 10^{3}\ \sqrt{0.01\over\epsilon}.$ (6)
For a fan beam the upper bound on $\theta$ may be as small as $\Omega$. Katz (2018, 2020) set other lower bounds on $\Gamma$ by requiring that it be large enough that the radiating charge bunches are not disrupted by their Coulomb repulsion.
3 Emission Region
The radiated beam is broadened beyond the intrinsic radiation pattern of accelerated charges if it is emitted by charges with imperfectly collimated velocity vectors. For curvature radiation in a neutron star’s magnetosphere the dimension $\delta r$ of the radiating region $\delta r\lesssim\theta r$. The radiation energy contained within a region of volume $(\delta r)^{3}$ is $E_{rad}=P_{rad}(\delta r/c)$ (this is much less that the energy of the burst if its width exceeds $\delta r/c\sim 10\,$ns). The energy density
${\cal E}_{rad}={E_{rad}\over(\delta r)^{3}}\sim 2\times 10^{32}\left({\epsilon\over 0.01}\right)^{-1/2}\left({I_{45}\over F_{Jy}}\right)^{-1/2}\ \text{erg/cm}^{3}.$ (7)
If the factors in parentheses are unity, ${\cal E}_{rad}$ equals the stress of a magnetic field of $7\times 10^{16}$ Gauss, and cannot be contained by plausible magnetar fields.
A small subset of magnetic field lines are not (in a magnetostatic approximation good to a fractional accuracy $\sim\omega r/c\sim 1\times 10^{-4}$) bent: for example, those along the axis of a dipole field. Along that axis the radiation may be emitted by region of length $\sim 0.3r$, multiplying the result of Eq. 7 by a factor $\delta r/(0.3r)\sim 10^{-3}$, reducing the corresponding magnetic field to $\sim 2\times 10^{15}$ Gauss, consistent with the polar field of a magnetar dipole and the spindown rate of FRB 20201124A. Higher multipole fields may be larger because they are effectively unconstrained by the spindown rate.
4 Discussion
Comparing the spindown rate and flux of FRB 20201124A estimates the collimation of its radiation and bounds the Lorentz factor of its emitting charges. It also suggests that the radiation is emitted along open field lines because their large radii of curvature imply a comparatively large volume can contribute to collimated radiation, relaxing the constraint on the magnetic field that confines the radiating charges. Alignment of the magnetic dipole and rotation axes of repeating FRB would explain the difficulty of finding their rotational periodicity because radiation from a field structure azimuthally symmetric around the rotation axis would not be rotationally modulated. It is implausible that neutron star fields are accurately dipolar near the surface, so this argument suggests that FRB are emitted at many neutron star radii. Emission along open field lines is consistent with, but does not require, that the periodic activity of some repeating FRB results from free precession of the neutron star (Zanazzi & Lai, 2020).
FRB emission along the magnetic dipole axis may also explain the distinction between repeating and non-repeating FRB: If aligned magnetic and rotational axes point in our direction, we may observe all the source’s bursts as a repeating FRB with intrinsic duty factor DF. If the magnetic and rotational axes are misaligned we can only observe bursts when, fortuitously and infrequently, the magnetic axis points toward us. We can observe only a fraction ${\cal O}(\Omega/4\pi)$ of all active repeaters close and powerful enough for detection because the others are beamed away from us. We can observe a large fraction of sufficiently close and powerful mis-aligned sources, but only as apparent non-repeaters with duty factors ${\cal O}(\text{DF}\Omega/4\pi)$; approximating the magnetic axes of the apparent non-repeaters as uniformly distributed on the sky, $\Omega/4\pi$ is the fraction of the time they are pointed in our direction. This is consistent with constraints (Katz, 2024b) on their duty factors, but both $\Omega$ and DF are uncertain even to order of magnitude.
If the preceding is valid, then the ratio in the local Universe of aligned to mis-aligned rotators is roughly the ratio of repeaters to non-repeaters in observations (such as those of CHIME-FRB) with a wide acceptance angle, or uncorrelated with the directions of known repeaters, provided the cumulative duration of the observation exceeds the intervals between bursts of repeaters. In practice, this is complicated by the facts that these intervals depend on observing sensitivity, likely vary among repeating sources and that the distribution of intervals between bursts is not Poissonian (Katz, 2024c).
The phenomenological differences (Pleunis et al., 2021) between bursts from repeaters and apparent non-repeaters, such as the “sad trombone” effect, must then be attributed to differences in the electrodynamics of aligned and mis-aligned rotators. This is consistent with the assumption that FRB are powered, like pulsars, by spin-down energy, not magnetostatic energy.
Data Availability
This theoretical study produced no new data.